7g of Compound X with molecular formula C4H8 are burned in a constant pressure carlomieter containing 25.00 kg of water at 25 C. The temp of the water is 3.011 C.

Calculate the standard hear formation of Compound X at 25 degrees C.
1
Expert's answer
2019-04-17T04:18:55-0400

All the heat released by the reaction is absorbed by the water:


"q_{absorbed} = cm\Delta T = 4186 \frac{J}{kg\times ^\circ C}\times 25.00 kg \times 3.011 ^\circ C= 315101 J = 315.1 kJ"

"\Delta H_{rxn} = -q_{absorbed}"

"\Delta H_{rxn} = -315.1 kJ"

"n(C_4H_8) = \frac{m(C_4H_8)}{M(C_4H_8)} = \frac{7 g}{56.11\frac{g}{mol}} = 0.1248 mol"

"\Delta H_f^0 = \frac{\Delta H_{rxn}}{n} = \frac{-315.1 kJ}{0.1248 mol} = -2525\frac{kJ}{mol}"